OpenStudy (anonymous):

the formula for the area of a trapezoid is 1/2 (a+b)h. apply the distributive property twice to write this formula in another way

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{2}(a+b)h\] \[\Large \frac{1}{2}(ah+bh) \ \text{Distribute the h through}\] \[\Large \frac{1}{2}ah+\frac{1}{2}bh \ \text{Distribute the 1/2 through}\] \[\Large \frac{ah}{2}+\frac{bh}{2} \ \text{Multiply}\] So either \(\Large \frac{1}{2}ah+\frac{1}{2}bh\) or \(\Large \frac{ah}{2}+\frac{bh}{2}\) work as an equivalent expression.

OpenStudy (anonymous):

thankyou ! makes sense !

jimthompson5910 (jim_thompson5910):

you're welcome